# Harnessing Large Language Models for Disaster Management: A Survey

### v1

@@ Abstract
Large language models (LLMs) have revolutionized scientific research with their exceptional capabilities and transformed various fields. Among their practical applications, LLMs have been playing a crucial role in mitigating threats to human life, infrastructure, and the environment. Despite growing research in disaster LLMs, there remains a lack of systematic review and in-depth analysis of LLMs for natural disaster management. To address the gap, this paper presents a comprehensive survey of existing LLMs in natural disaster management, along with a taxonomy that categorizes existing works based on disaster phases and application scenarios. By collecting public datasets and identifying key challenges and opportunities, this study aims to guide the professional community in developing advanced LLMs for disaster management to enhance resilience against natural disasters.
@@
Zhenyu Lei   Yushun Dong  Weiyu Li Rong Ding   Qi Wang   Jundong Li University of Virginia, Florida State University, Northeastern University {vjd5zr, jundong}@virginia.edu, yd24f@fsu.edu {weiy.li, ding.ro, q.wang}@northeastern.edu
@@ 1 Introduction
Natural disasters have become increasingly frequent and severe, posing unprecedented threats to human life, infrastructure, and the environment Manyena (2006); Yu et al. (2018); Chaudhary and Piracha (2021). The 2010 Haiti earthquake, for instance, led to more than 200,000 fatalities and widespread infrastructure devastation DesRoches et al. (2011). Similarly, the 2020 Australian bushfires have resulted in the deaths of at least 33 people and an estimated one billion animals Deb et al. (2020). The profound impacts of such catastrophic events have highlighted the urgent need for effective disaster management strategies. Recently, large language models (LLMs) have revolutionized the scientific research area due to their unparalleled abilities in contextual understanding, logical generation, and complex problem-solving with textual data and data from various modalities Zhang et al. (2024b, a). These capabilities make LLMs well-suited for natural disaster management, with models capable of analyzing vast disaster real-time data, facilitating dynamic communication with affected communities, and supporting critical decision-making Otal et al. (2024).
Despite the potential of LLMs in disaster management, a systematic review is still missing. This absence hinders researchers and practitioners from identify best practices, uncovering research gaps, and optimizing the deployment of LLMs for effectively addressing disaster-related challenges. To address this gap, this paper systematically surveys the applications of LLMs based on three types of architectures across the four disaster phases: mitigation, preparedness, response, and recovery. We then propose a novel taxonomy that integrates application scenarios, specific tasks, and the architectures of models addressing these tasks. Furthermore, we summarize publicly available datasets, identify key challenges, and investigate avenues for improving the effectiveness, efficiency, and trustworthiness of LLMs. This review aims to inspire and guide AI researchers, policymakers, and practitioners towards developing LLM-aided disaster management frameworks. Our contributions are listed below:
  • We systematically introduce the current explorations handling disaster management with LLMs across four disaster phases.
  • We propose a taxonomy integrating application scenarios, tasks, and model architectures, offering practical and technical insights.
  • We collect key resources such as public datasets and highlight the current challenges and future research opportunities.
Figure 1: Taxonomy of applications of LLMs in disaster management. This survey categorizes the utilization of LLMs across four disaster phases, highlighting specific applications where three types of LLMs (Encoder-based, Decoder-based, and Multimodal LLM) perform tasks such as classification, estimation, extraction, and generation. The chart in the bottom-right corner presents the distribution of surveyed papers across each phase.
@@ 2 Background
Disaster management is a multidisciplinary field focused on coordinating resources, expertise, and strategies to address the challenges posed by increasingly severe disasters, minimizing immediate damage while fostering long-term resilience and adaptive recovery. Disaster management consists of four interconnected phases Sun et al. (2020):
  • Mitigation. This phase involves identifying risks and vulnerabilities and implementing proactive measures for disaster prevention.
  • Preparedness. This phase includes comprehensive planning and public education to ensure readiness for upcoming disasters.
  • Response. This phase identifies and addresses immediate needs during a disaster, including rescue and resource distribution.
  • Recovery. This phase involves rebuilding affected areas, addressing both physical and social impacts to facilitate a return to normalcy.
In general, LLMs are able to serve as a general-purpose foundation for building specialized AI tools in disaster management. Here, we categorize LLM architectures into: (1) encoder-based LLM (e.g., BERT Devlin (2018)) primarily for extracting textual features for downstream tasks; (2) decoder-based LLM (e.g., GPT Brown (2020)) for text generation from given textual queries; (3) multimodal LLMs that integrate multiple modalities via alignment Tiong et al. (2022) or feature fusion Madichetty et al. (2021). For disaster management, downstream tasks usually include classification (e.g., disaster detection), estimation (e.g., severity estimation), extraction (e.g., knowledge extraction), and generation (e.g., report generation). Techniques such as fine-tuning, prompting, and additional adapters are commonly employed to adapt LLMs for specific downstream applications.
@@ 3 LLM For Disaster Management
Foundation models can be applied to four phases of disaster management, namely mitigation, preparedness, response, and recovery. For each phase, works are organized by application scenarios, tasks, and model architectures. We present the taxonomy in Figure 1 and summaries in Appendix A.
@@ 3.1 Disaster Mitigation
Assessing vulnerabilities is a critical aspect of disaster mitigation, where LLMs have shown promising potential. This process typically involves identifying and analyzing infrastructure or communities that are susceptible to disasters.
Vulnerability Classification. Infrastructure Ombudsman has compared the supervised classification with encoder-based LLMs and zero-shot prompt learning with decoder-based LLMs in detecting specific types of concerns about potential infrastructure failures from social media data Chowdhury et al. (2024), facilitating decision-makers in effectively prioritizing resources.
Answer Generation. Apart from the infrastructure vulnerability, decoder-based LLMs can assist in answering community vulnerability-related queries based on retrieval of social vulnerability index (SVI) Martelo and Wang (2024).
@@ 3.2 Disaster Preparedness
In the long term, LLMs can play a pivotal role in disaster preparation via (1) public awareness enhancement and consistent (2) disaster forecast. Building on the forecasts, LLMs can support decision-makers in issuing (3) disaster warnings to the public, which can enhance preparedness in the short term. Furthermore, LLMs can assist in well-structured (4) evacuation planning, ensuring safe relocation of individuals and property.
@@ 3.2.1 Public Awareness Enhancement
It is critical to enhance public awareness regarding disasters by providing insights and knowledge extracted from previous disaster experiences.
Knowledge Extraction. Encoder-based LLMs are particularly valuable due to their exceptional knowledge extraction capabilities. BERT has been fine-tuned to extract disaster-related knowledge from news and social media Fu et al. (2024), and extensive literature Zhang and Wang (2023). To enhance the logical coherence of extracted entities, Ma et al. propose BERT-BiGRU-CRF for name entity and relation recognition (NER) to construct disaster knowledge graphs Ma et al. (2023). With enhanced reasoning ability, decoder-based LLMs have been utilized, where ChatGLM2 is fine-tuned to extract knowledge triplets from documents to construct a knowledge graph Wu et al. (2024).
Answer Generation. The derived knowledge could be incorporated in decoder-based LLMs’ prompts, which could facilitate answering wildfire Hostetter et al. (2024) and flood risk-related questions Martelo and Wang (2024), and offering personalized mitigation guidance Li et al. (2023). In addition, techniques such as retrieval-augmented generation (RAG) have been utilized for knowledge incorporation, with knowledge retrieved from flood and geographic information system (GIS) knowledge graph Zhu et al. (2024).
3.2.2 Disaster Forecast
Effective disaster preparedness also hinges on accurate and robust forecasting and risk level prediction.
Occurrence Classification. With their extensive internal knowledge, encoder-based LLMs have proven to be powerful tools for classification-based disaster forecasting. For instance, Indra et al. streamline BERT, GRU, and CNN to anticipate disaster Indra and Duraipandian (2023), improving their ability to capture temporal dependencies and local patterns in large-scale social media text. However, textual data is limited due to its subjective and imprecise nature, which leads to multimodal LLMs integrating various data modalities. Zeng et al. combine historical flood data and BERT embeddings extracted from geographical description, then utilize XGBoost to predict the disaster risk Zeng and Bertsimas (2023). Debata et al. further extract features from statistical data with k-nearest neighbors (KNN) for better incorporation. Further advancements incorporate additional satellite imagery features derived from ResNet for more informative disaster prediction Liu and Zhong (2023). To facilitate disaster risk prediction with explicit external knowledge, decoder-based LLMs have been utilized with RAG to retrieve historical flood data from disaster knowledge base to derive risk level and action recommendation Wang et al. (2024).
@@ 3.2.3 Disaster Warning
Once a disaster has been foreseen, timely warnings and suggestions are crucial to public safety.
Warning Generation. Decoder-based LLMs have been a valuable tool in generating clear warning messages from rule-based alerts derived from streaming data Chandra et al. (2024), significantly enhancing the responsiveness of warning systems. In addition, LLMs have been utilized to retrieve disaster alerts from National Weather Service’s API to provide detailed, real-time information about upcoming disaster Martelo and Wang (2024).
Image Generation In addition to textual warnings, visual warnings can convey vivid descriptions, reaching a broader audience. To achieve this, multimodal LLMs enhanced by diffusion-based text-to-image generative models can produce vivid visual representations of impending disasters Lubin et al. (2024), improving the clarity of alerts.
@@ 3.2.4 Evacuation Planning
Plan Generation
To ensure the safety of individuals and properties, structured evacuation plans have been developed. Leveraging the exceptional reasoning capabilities, decoder-based LLMs have been prompted to plan escapes and provide evacuation suggestions Hostetter et al. (2024), thus enhancing public safety against incoming disasters.
@@ 3.3 Disaster Response
With accurate and real-time (1) disaster identification and (2) situation assessment, decision makers can gain essential knowledge to form a response basis. In addition, LLM can facilitate the (3) disaster information coordination, thus boosting a collaborative disaster response among stakeholders. Consequently, decision makers can execute effective actions empowered by LLMs, such as (4) disaster rescue, and offering (5) disaster issue consultation.
@@ 3.3.1 Disaster Identification
Effective disaster response begins with accurate and real-time identification to enable efficient interventions Said et al. (2019); Weber et al. (2020), with social media serving as a valuable avenue by providing real-time updates from affected individuals Anderson (2016); Trono et al. (2015).
Relevance Classification. Leveraging their advanced capabilities, LLMs have become prominent disaster detection tools, particularly by performing binary classification for disaster-related posts.
Encoder-based LLMs, with a trainable adapter appended, are widely utilized in this application via fine-tuning Ningsih and Hadiana (2021); Singh et al. (2022). Beyond fine-tuning existing LLMs, researchers have developed novel techniques tailored to specific disaster contexts. CrisisTransformers has pre-trained transformers initialized randomly or with RoBERTa weights on a large dataset with millions of crisis-related tweets before fine-tuning for specific downstream tasks Lamsal et al. (2024a), thus enhancing extensive disaster knowledge incorporation. Additionally, Paul et al. tackle the challenge of limited labeled data by incorporating active learning, which fine-tunes transformer-based representations to prioritize the labeling of the most informative samples Paul et al. (2023).
More innovations focus on framework design, where ensemble models have been employed to integrate the strengths of various LLMs Mukhtiar et al. (2023). Furthermore, additive modifications also enhance LLMs. For instance, several studies append CNN and att-BiLSTM to capture temporal dependencies and local patterns Manthena (2023). In addition, graph neural networks (GNN) can enhance word relation modeling. For instance, Danday et al. construct a word relationship graph from DistilBERT embeddings, which are sent to GNN for classification Danday and Murthy (2022). For better modality fusion, GNom integrates graph features derived from multilingual word graphs and BERT text embeddings with a novel cross-attention mechanism Ghosh et al. (2022).
In addition, advanced decoder-based LLMs such as Llama-2 and GPT-4 demonstrate exceptional contextual understanding ability, being prompted for classification Taghian Dinani et al. (2023).
Image data also offer valuable insights, where multimodal LLMs—combining text and other images—has emerged as a promising approach, enhancing the accuracy of disaster identification. Several approaches combine text features from encoder-based LLMs and image features by aggregation Kamoji et al. (2023); Madichetty et al. (2021). In addition, Koshy et al. append a BiLSTM module after RoBERTa for better sequential modeling Koshy and Elango (2023). To improve modality alignment, Shetty et al. utilize the attention mechanism to fuse text and image features Shetty et al. (2024), while Zhou et al. adopt a Cycle-GAN with mixed fusion strategy to address the multimodal heterogeneity challenges Zhou et al. (2023b). Beyond modality alignment, aligning features across various disaster events is critical in label-scarce scenarios. Jang et al. introduce a semi-supervised minimax entropy domain adaptation framework unifying multimodal features from various disaster events, enabling generalization across diverse contexts Wang and Wang (2022). To combine strengths of diverse LLMs and visual models, the ensemble is further utilized Hanif et al. (2023).
In addition to social media data, other real-time data sources, such as satellite imagery and news articles, can also provide valuable insights. For instance, Jang et al. utilize ResNet to process satellite images and BERT to process descriptions of the disasters from news Jang et al. (2024).
@@ 3.3.2 Disaster Situation Assessment
After identifying a disaster, evaluating its severity and spread is crucial for response strategies. This information is often embedded in textual data, such as social media posts reporting situation Saroj and Pal (2020), and news reports offering precise descriptions Houston et al. (2012); Sood et al. (1987).
Situation Classification. Madichetty et al. fine-tune encoder-based LLMs to identify situational posts Madichetty and Sridevi (2021), combining the RoBERTa and SVM representations for binary classification. Raj et al. utilize BERT to filter disaster-related tweets, followed by NER to extract locations, with location counts indicating disaster severity Raj et al. (2023). In addition, visual data from the web and satellites is also integrated and processed by multimodal LLMs to enhance the disaster assessment. Kanth et al. utilize BERT and CNN to filter disaster-related tweets and classify the disaster severity based on the images associated with informative tweets Kanth et al. (2022).
Severity Estimation. Classification only offers a coarse understanding, while severity estimation provides precise quantitative insights. Mousavi et al. utilize chain-of-thoughts (CoT) prompting on decoder-based LLMs (e.g. Gemini) to estimate the earthquake intensity as Modified Mercalli Intensity (MMI) values Mousavi et al. (2024). In addition, multimodal LLMs can be utilized to perform image-related estimation. For instance, FloodDepth-GPT uses prompts to guide GPT-4 to estimate the floodwater depth of the floodwater Akinboyewa et al. (2024) from flood images and descriptions.
Description Generation. Beyond categorical and statistical descriptions, multi-modal LLMs can generate more comprehensible textual situational descriptions. For instance, GPT-4V is prompted to translate situational images to descriptions Hu and Rahnemoonfar (2024), while Wolf et al. prompts GPT-4 to generate assessment descriptions Wolf et al. (2023) with disaster image data.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains, including their emerging role in mitigating threats to human life, infrastructure, and the environment during natural disasters. Despite increasing research on disaster-focused LLMs, there remains a lack of systematic reviews and in-depth analyses of their applications in natural disaster management. To address this gap, this paper presents a comprehensive survey of LLMs in disaster response, introducing a taxonomy that categorizes existing works based on disaster phases and application scenarios. By compiling public datasets and identifying key challenges and opportunities, this study aims to provide valuable insights for the research community and practitioners in developing advanced LLM-driven solutions to enhance resilience against natural disasters.
@@
Zhenyu Lei   Yushun Dongthanks: Corresponding authors  Weiyu Li Rong Ding   Qi Wang11footnotemark: 1   Jundong Li11footnotemark: 1 University of Virginia, Florida State University, Northeastern University {vjd5zr, jundong}@virginia.edu, yd24f@fsu.edu {weiy.li, ding.ro, q.wang}@northeastern.edu
@@ 1 Introduction
Natural disasters are becoming increasingly frequent and severe, posing unprecedented threats to human life, infrastructure, and the environment Manyena (2006); Yu et al. (2018); Chaudhary and Piracha (2021). The 2010 Haiti earthquake, for instance, resulted in over 200,000 fatalities and widespread infrastructure devastation DesRoches et al. (2011). Similarly, the 2020 Australian bushfires caused the deaths of at least 33 people and an estimated loss of one billion animals Deb et al. (2020). The profound impact of such catastrophic events underscores the urgent need for effective disaster management strategies. Recently, large language models (LLMs) have transformed research and technological innovation with their exceptional capabilities in contextual understanding, logical reasoning, and complex problem-solving across multiple modalities Zhang et al. (2024b, a); Wei and Liu (2025). These capabilities position LLMs as powerful tools for natural disaster management, enabling them to analyze vast real-time disaster data, facilitate dynamic communication with affected communities, and support critical decision-making Otal et al. (2024).
Despite their potential, a systematic review of LLMs in disaster management remains absent, limiting researchers and practitioners in identifying best practices, addressing research gaps, and optimizing LLM deployment for disaster-related challenges. To bridge this gap, this paper presents a comprehensive survey of LLM applications in disaster management, categorizing them across three model architectures and the four key disaster phases: mitigation, preparedness, response, and recovery. We introduce a novel taxonomy that integrates application scenarios, specific tasks, and model architectures tailored to disaster-related challenges. Additionally, we summarize publicly available datasets, identify key challenges, and explore avenues for enhancing the effectiveness, efficiency, and trustworthiness of LLMs in disaster response. This review aims to inspire and guide AI researchers, policymakers, and practitioners toward developing LLM-driven disaster management frameworks. Our key contributions are as follows:
  • Systematical Review: We provide the first systematical review of explorations of LLMs applications in disaster management across four key disaster phases.
  • Novel Taxonomy: We propose a taxonomy integrating application scenarios, specific tasks, and model architectures, providing both practical and technical insights into this survey.
  • Resource Compilation: We compile essential resources (e.g., datasets), and highlight key challenges and future research directions to advance LLM-driven disaster management.
@@ 2 Background
Disaster management is a multidisciplinary field that integrates resources, expertise, and strategies to mitigate the impact of increasingly severe disasters. Its primary goal is to minimize immediate damage while fostering long-term resilience and adaptive recovery. Disaster management comprises four interconnected phases Sun et al. (2020):
  • Mitigation involves identifying risks and vulnerabilities while implementing proactive measures to prevent disasters.
  • Preparedness includes developing comprehensive plans and public education initiatives to enhance readiness for potential disasters.
  • Response identifies and addresses immediate needs during a disaster, including emergency rescue operations and resource distribution.
  • Recovery involves rebuilding affected areas, addressing both physical and social impacts to facilitate a return to normalcy.
In general, LLMs have the potential to serve as general-purpose foundations for developing specialized AI tools that enhance various aspects of disaster management. Here, we categorize LLM architectures into three main types: (1) encoder-based LLM (e.g., BERT Devlin (2018)), which excel in contextual understanding; (2) (encoder-)decoder LLM (e.g., GPT Brown (2020)), which are optimized for sequential prediction; and (3) multimodal LLMs, which integrate multiple modalities to enhance information processing Tiong et al. (2022); Madichetty et al. (2021) In disaster management, common downstream tasks include classification (e.g., damage classification), estimation (e.g., severity estimation), extraction (e.g., knowledge extraction), and generation (e.g., report generation). To tailor LLMs for these tasks, techniques such as fine-tuning and prompting are commonly employed.
@@ 3 LLM For Disaster Management
Foundation models can be utilized across the four disaster management phases: mitigation, preparedness, response, and recovery. Within each phase, existing works are categorized based on application scenarios, specific tasks, and model architectures. Figure 1 presents an overview of our taxonomy, with detailed summaries provided in Appendix A.
@@ 3.1 Disaster Mitigation
Assessing vulnerabilities is a crucial component of disaster mitigation, where LLMs have demonstrated promising potential. This process involves identifying and analyzing infrastructure and communities at risk, enabling proactive measures to reduce disaster impact.
Vulnerability Classification. A system named Infrastructure Ombudsman has leveraged supervised learning with encoder-based LLMs and zero-shot prompt learning with (encoder-)decoder LLMs to detect and classify concerns about potential infrastructure failures from social media data Chowdhury et al. (2024). This approach enables decision-makers to effectively prioritize resources and address critical issues in a timely manner.
Answer Generation. Beyond infrastructure vulnerability assessment, (encoder-)decoder LLMs can help address community vulnerability-related queries by retrieving and leveraging the Social Vulnerability Index (SVI) Martelo and Wang (2024).
@@ 3.2 Disaster Preparedness
In the long term, LLMs can play a pivotal role in disaster preparedness through (1) enhancing public awareness by disseminating accurate and accessible information, and (2) supporting disaster forecasting with advanced data analysis. Building on these forecasts, LLMs can aid decision-makers in issuing (3) timely disaster warnings, improving short-term preparedness. Furthermore, LLMs can support well-structured (4) evacuation planning, ensuring the safe relocation of individuals and assets.
Figure 1: Taxonomy of applications of LLMs in disaster management. This survey categorizes the utilization of LLMs across four disaster phases, highlighting specific applications where tasks such as classification, estimation, extraction, and generation are performed by three types of LLMs (Encoder-based, (Encoder-)Decoder, and Multimodal LLM). The chart in the bottom-right corner presents the distribution of papers across each phase.
@@ 3.2.1 Public Awareness Enhancement
Enhancing public awareness of disasters is crucial, particularly by providing insights and knowledge derived from past disaster experiences.
Knowledge Extraction. Encoder-based LLMs have been fine-tuned to extract disaster-related knowledge from news articles and social media Fu et al. (2024), as well as from extensive disaster literature Zhang and Wang (2023), using Named Entity Recognition (NER). To improve the logical coherence of extracted entities, Ma et al. propose BERT-BiGRU-CRF for NER, enabling the construction of disaster knowledge graphs Ma et al. (2023). In addition, (encoder-)decoder LLMs have been fine-tuned with instructional learning to extract knowledge triplets from documents for knowledge graph construction Wu et al. (2024).
Answer Generation. The extracted disaster knowledge could be incorporated in (encoder-)decoder LLMs’ prompts, facilitating disaster-related question answering Hostetter et al. (2024); Martelo and Wang (2024); Li et al. (2023). Additionally, techniques such as retrieval-augmented generation (RAG) have been employed to further improve knowledge integration Zhu et al. (2024).
3.2.2 Disaster Prediction
Effective disaster preparedness also relies on accurate and reliable disaster prediction.
Occurrence Classification. Encoder-based LLMs have been widely employed for disaster prediction. For instance, BERT has been integrated with GRU and CNN to predict disasters Indra and Duraipandian (2023). However, textual data alone is often limited due to its subjective and imprecise nature, prompting the adoption of multimodal LLMs that incorporate multiple data modalities. For instance, Zeng et al. combine historical flood data with geographical descriptions of specific locations to assess disaster risk Zeng and Bertsimas (2023). Additionally, satellite imagery has been leveraged to provide visual context, enhancing predictive accuracy Liu and Zhong (2023). To further improve disaster prediction with explicit external knowledge, (encoder-)decoder LLMs have been integrated with RAG to retrieve historical flood data, aiding in risk assessment and action recommendation Wang et al. (2024).
@@ 3.2.3 Disaster Warning
Once a disaster is anticipated, timely warnings are essential for ensuring public safety.
Warning Generation. (Encoder-)decoder LLMs have proven valuable in generating warning messages based on rule-based alerts derived from streaming data Chandra et al. (2024), significantly improving the responsiveness of warning systems. Additionally, RAG has enhanced LLMs by enabling the retrieval of disaster alerts from official APIs, providing real-time information on impending disasters Martelo and Wang (2024).
Image Generation. In addition to textual warnings, visual warnings can provide more vivid and intuitive descriptions, effectively reaching a broader audience. To achieve this, multimodal LLMs enhanced by diffusion-based text-to-image generative models can generate detailed visual representations of impending disasters Lubin et al. (2024), enhancing the clarity and impact of disaster alerts.
@@ 3.2.4 Evacuation Planning
Plan Generation.
To safeguard individuals and property from impending disasters, (encoder-)decoder LLMs have been prompted to generate escape plans and provide evacuation recommendations Hostetter et al. (2024).
@@ 3.3 Disaster Response
With accurate and real-time (1) disaster identification and (2) situation assessment, decision-makers can acquire critical insights to establish a solid foundation for response efforts. Additionally, LLMs can facilitate (3) disaster information coordination, enhancing collaboration among stakeholders for more effective disaster response. As a result, decision-makers can leverage LLMs to execute key actions, including (4) disaster rescue operations and (5) disaster-related consultations.
@@ 3.3.1 Disaster Identification
Effective disaster response begins with accurate and real-time identification, enabling efficient interventions Said et al. (2019); Weber et al. (2020). Social media serves as a valuable resource in this process, offering real-time updates from affected individuals Anderson (2016); Trono et al. (2015).
Relevance Classification with Encoder-based LLMs. Classifying social media posts to identify disaster-related content is a crucial step in disaster detection, where LLMs have proven to be highly effective. Encoder-based LLMs augmented with trainable adapters are commonly employed for this task through fine-tuning on annotated disaster corpora Ningsih and Hadiana (2021); Singh et al. (2022); Lamsal et al. (2024a). Recognizing the diverse sources of disaster data, ensemble methods combine predictions from multiple LLMs to leverage their complementary strengths in processing varied linguistic patterns Mukhtiar et al. (2023). Pure LLM-based approaches may struggle to capture fine-grained structural features in disaster-related posts. To address this, hybrid architectures integrate CNNs to capture local n-gram patterns Franceschini et al. (2024); Song and Huang (2021); Meghatria et al. (2024), attention-based BiLSTMs to model sequential dependencies Huang et al. (2022), and graph neural networks (GNNs) to represent semantic word relationships Manthena (2023); Ghosh et al. (2022). To tackle the challenge of limited labeled training data, active learning has been employed to automatically label informative samples Paul et al. (2023).
Relevance Classification with Encoder-based LLMs. Furthermore, (encoder-)decoder LLMs such as GPT-4 have demonstrated strong performance in relevance classification using prompt learning techniques Taghian Dinani et al. (2023).
Relevance Classification with Multimodal LLMs. Image data also provide valuable insights for disaster analysis and can be integrated to enhance classification using multimodal LLMs. This integration can be achieved through simple aggregation Kamoji et al. (2023); Madichetty et al. (2021) or attention-based mechanisms Shetty et al. (2024). To address challenges arising from multimodal heterogeneity, Zhou et al. employ a Cycle-GAN combined with a mixed fusion strategy Zhou et al. (2023b). Beyond multimodal heterogeneity, research also tackles other challenges in multimodal learning. These include addressing label scarcity through semi-supervised minimax entropy domain adaptation frameworks Wang and Wang (2022) and enhancing model performance by leveraging the complementary strengths of diverse LLMs and visual models using ensemble methods Hanif et al. (2023). Beyond social media, data from sources such as satellite imagery and news articles can further enhance disaster analysis Jang et al. (2024).
@@ 3.3.2 Disaster Situation Assessment
After disaster identification, assessing its severity and spread is essential for formulating effective response strategies.
Situation Classification. encoder-based LLMs have been fine-tuned to for binary classification to identify situational posts Madichetty and Sridevi (2021). Raj et al. employ BERT and NER to extract disaster-related locations, using location counts as an indicator of disaster severity Raj et al. (2023). Additionally, multimodal LLMs integrate visual data to further enhance disaster situational assessment Kanth et al. (2022).
Severity Estimation. While classification provides only a coarse understanding, severity estimation offers precise quantitative insights. (encoder-)decoder LLMs enhanced with chain-of-thought (CoT) reasoning have been used to estimate earthquake intensity, expressed as Modified Mercalli Intensity (MMI) Mousavi et al. (2024). In addition, multimodal LLMs use rich image data for more accurate estimations. For example, FloodDepth-GPT employs prompt-based guidance with GPT-4 to estimate floodwater depth from flood images.
Description Generation. Beyond categorical and statistical descriptions, multimodal LLMs can generate more comprehensible textual situational reports from disaster images Hu and Rahnemoonfar (2024); Wolf et al. (2023).